geerky42 (geerky42):

A body moves with decreasing speed. Kinetic Energy is increasing or decreasing? What about net work? Why?

geerky42 (geerky42):

I believe net work is negative since there should be a force applying to body at opposite direction. Not sure about kinetic energy...

OpenStudy (anonymous):

@geerky42 KE is directly proportional to square of speed so if speed decreases K.E. is also decreasing while net work is done -ve or +ve depends on that whether you want to find work done by the force which makes it speed decreasing is -ve otherwise work done by the object whose speed is decreasing is +ve:)

OpenStudy (anonymous):

if veleocity decres its momentum n thus kinetc enrgy decreses..

OpenStudy (anonymous):

Kinetik energy is deacrease and : \[Work = \left( \frac{ 1 }{ 2 } \times m \times v _{2} ^{2} - \frac{ 1 }{ 2 } \times m \times v _{1} ^{2} \right)\] \[Work = \frac{ 1 }{ 2 } \times m \times \left( v _{2} ^{2} -v _{1} ^{2} \right) \]